# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

### 16 Dec 2023 v1

@@ Introduction
Concerning the decisions grounds, for almost every platform incompatible content was the overwhelming majority, up to 100% in the case of LinkedIn. The exception was X, which indicated illegal content for all of its SoRs. Only Snapchat indicated an overlap of incompatible and illegal content in 12.2% of their SoRs. None of the other social media platforms used this attribute.
Figure 1: Platform-wise distribution of infringement categories of statement of reasons (SoR), seriated using Manhattan distance across platforms and categories, so that similar platforms are positioned close to one another.
@@
Curiously, category addition was merely used by Pinterest and X, but in less than 0.04% of their SoRs. As for keywords, only half of the social media platforms used them for category specification: Pinterest, Snapchat, X, and YouTube. However, these are not always informative. For instance, circa 75% of YouTube’s SoRs have a keyword —all of them other— which YouTube explained by specifying that “[their] mapping has not yet been completed”. Only 10% of Snapchat’s SoRs have a keyword, 86% of these being sexual abuse material and the rest online bullying intimidation. About 87% of Pinterest’s SoRs have a keyword (with a few having two), predominantly adult sexual material. Interestingly, all of X’s SoRs have a keyword (with a few having two or three), but almost half (45%) is other (detailed as a policy violation), while 38% is adult sexual material. The rest is child sexual abuse material, nudity, and goods/services not permitted.
@@
[…]
Table 2: Platform-wise distribution of decision types, ordered alphabetically and by type mean across platforms.
Figure 2: Platform-wise distribution for decisions of the visibility type, seriated using Manhattan distance across platforms and sub-types, so that similar platforms are positioned close to one another.
Figure 3: Distribution of moderated content type by platform. Custom other names were simplified for space reasons.
Figure 4: Distribution of decision application delays (with respect to the date of content creation) by social media platform, ordered by their 80% winsorized means (W).
Types of Restriction Decisions
There are four main types of restriction decisions: account (suspension, termination), monetary (suspension, termination, other), provision (partial/total suspension/termination of the service), and visibility (restriction for the content). Platforms must indicate a single sub-type of one of the four types. However, these types are not mutually exclusive, meaning that a SoR can contain multiple types of restrictions, with at least one of them being mandatory. However, all of the social media platforms set a single type. As detailed in Table 2, most were in visibility (78.7% across platforms), followed by account (18.1%) and provision (3.1%). Surprisingly, there was not a single SoR with a monetary restriction decision. For visibility sub-types, shown in Figure 2, we noticed a similar platform grouping to that of infringement categories:
  1. 1.
@@
The required attribute content type is a list with one or more of the following predefined values: app, audio, image, product, synthetic media (e.g., content generated via AI), text, video, and other. Despite the possibility to indicate multiple values, all eight platforms only specified a single value for all of their SoRs, except YouTube, which however indicated multiple values only in a handful of cases (<<<0.01%). We found that social media platforms greatly differ in how they specify the content type of their SoRs, as shown on the left side of Figure 3. Pinterest and LinkedIn only set content type other, and did not map any of their internal types to one or more of the predefined values. YouTube also had a remarkable share of other content (91.1%), prevalently ads. Facebook and Instagram have similar distributions, with most content being other (almost 60%), all detailed as accounts, which is consistent with their distributions of decision types in Table 2. Incidentally, Facebook was the only social media with content type product (1.2%). Most of Snapchat’s moderated content was video, with other content (24%) being mainly accounts. Still, 16.6% of other was labelled as multi-media, instead of using multiple predefined values as prescribed by the DSA-TDB. Interestingly, TikTok’s moderated content was mostly text (55.5%) and not video (33.8%), with a small share of other content (1.3%), primarily accounts. Unexpectedly, moderated content by X was overwhelmingly synthetic media (99.8%), with the second being audio (0.01%) —the only social media to use this type— and only a handful of text.
Timeliness
@@
Artice 17 of the DSA requires that platforms communicate their SoRs without undue delay to allow close to real-time updates. Pinterest, Snapchat, and TikTok started submitting SoRs on the very day that the DSA-TDB was made available (i.e., on 25 September). Facebook, Instagram, YouTube, and LinkedIn the following day. X was last in line and started a week later (i.e., on 3 October).
Figure 5: Platform-wise distribution of automated infringement detection and decision, ordered by an ad hoc automation index.
@@
We then analyzed the delay with which social media communicated SoRs, measured as the difference in days between the date of application of the moderation intervention and the date of communication in the corresponding SoR (see Figure 8 in the Appendix). LinkedIn, Pinterest, and Snapchat for the most part communicated their SoRs on the same day of application and with little to no fluctuations. Despite the late start, X caught up on the first week of submissions by communicating content moderated since 25 September. Afterwards, it began to communicate SoRs the same day of their application. On the contrary, we noticed that TikTok started by stably submitting on the same day of application, but at the end of the 60-day period it communicated SoRs with an uprising mean delay of up to six days. On the other hand, Facebook, Instagram, and YouTube communicated their SoRs with a mean delay of one day, but with marked fluctuations throughout the period.
@@
Arguably, a prompt response by platforms in moderating an infringing content is desirable. Hence, we also measured the delay in days of the restriction decision with respect to the creation of the infringing content. We observed great heterogeneity in the distributions of these delays among platforms. For this reason, in their analysis and visualization we opted for using few categorical cuts, as shown in Figure 4. In addition, we also report in the figure the 80% winsorized mean (W) as a robust measure of centrality that lessens the influence of outliers. Oddly, X was the only platform that always applied a decision on the same day the infringing content was created (W===0). TikTok had the second lowest delay, with 89.7% of decisions being applied the day of content creation. Next, LinkedIn applied its decision within 7 days in 89.6% of their SoRs, while Instagram in 81.4%. Even if Snapchat had a higher value for the same timeframe (82.5%), in 14.1% of SoRs it took more than 30 days to apply its decision. Facebook presented a similar but even more marked situation, with 84.2% and 22.5%, respectively. YouTube, on the other hand, had a substantial amount of decisions that took more than 30 days (68.4%), while also having many within 15–30 days (15.3%). However, the distribution with the highest mean and dispersion was that of Pinterest, as 16.9% of decisions were applied the same day of content creation while 71.6% after 30 days, with many even taking hundreds of days.
[…]
Automated Means
@@
Automation can be applied at various steps in the content moderation process. It is often used to scale up the detection of offending content and to speed up the decision on how to moderate it. Consequently, the DSA-TDB includes two attributes to describe the possible use of automation for moderation: automated detection (yes, no) and automated decision (not, partially, fully). Figure 5 provides a combined view of the use of automation for detection and decision by social media platforms. Again, we see different approaches being adopted. Surprisingly, none of the SoRs submitted by X specified the use of automated means whatsoever. Snapchat is second in terms of non-automated means in their SoRs (76%). Pinterest mainly uses non-automated detection with partially automated decision (98%), whereas LinkedIn mainly uses automated detection together with non-automated decision (68%). YouTube is the only platform that uses all combinations of automated means, with automated detection and non-automated decision being most common (38%), followed by fully automated means (31%), and non-automated means (28%). Once more, Instagram and Facebook have similar distributions, with automated detection and partially automated decision being the most common for both, with 92% and 99%, respectively. Almost opposite of X, TikTok predominantly indicates fully automated means for moderation (92%).
@@
To provide a holistic view of social media behavior with respect to automated moderation, we converted the boolean and ordinal attributes of the DSA-TDB into two numeric indices of automated detection and automated decision. Both indices are defined in the [0, 1] range and are valued, for each platform, depending on the frequency with which automated detection and decision occurred in the platform’s SoRs. We were then able to position each platform in the bi-dimensional space depicted in Figure 6, where dot size is proportional to the number of SoRs per MAAR, and dot color represents the mean of the automated detection and automated decision indices. The platform positioning within the newly defined automation space reveals four groups of platforms with overall similar degrees of automation: low (X and Snapchat), modest (Pinterest), substantial (LinkedIn, YouTube, Instagram, and Facebook), and high (TikTok).
Comparison with Transparency Reports
@@
Within the DSA framework there are multiple mechanisms to ensure transparency in content moderation, including the DSA-TDB that we analyzed so far, and the DSA Transparency Reports. Both resources cover similar aspects of the content moderation process, albeit at different granularity. As such, there is overlapping between the information contained in the two resources, which opens up the possibility to carry out thorough comparisons. In this section we leverage this opportunity to carry out the first systematic comparison between the data submitted by each social media to the DSA-TDB and that declared in the corresponding transparency report. In detail, for each platform and for each dimension of moderation (e.g., timeliness, automation, etc.), we look for the corresponding information in that platform’s most recent transparency report. We then compare the two figures to corroborate or challenge the self-declared information provided by each platform. This comparison serves as a validation step of the DSA-TDB data, and allows for spotting inconsistencies and discrepancies. In general, we were not able to compare each dimension of moderation for each of the eight platforms, due to differences in the data they provided in their transparency reports. However, all platforms —with the exception of Snapchat— provided detailed information about the use of automated means for moderation, for which we found many inconsistencies. In the case of X, however, we found inconsistencies in multiple aspects. In the following, we present the results of our comparisons starting from the platforms for which we found less issues.
@@
Pinterest, Facebook, Instagram, LinkedIn. Pinterest’s self-reporting was overall congruous and detailed. Interestingly, it was the only social media to use three automation levels in their transparency report —manual, hybrid, and automated— similarly to the organization of the DSA-TDB. Unsurprisingly, Facebook and Instagram share similar transparency reports, based on Meta guidelines. In their reports, both state the heavy use of automated means, including full automation for both for detection and decision. However, none of their SoRs specified fully automated decisions, let alone full automation overall, as shown in Figure 5. For the reporting period 25 April – 30 September (159 days), Meta platforms automatically removed the following organic content (i.e., non-paid): for Facebook, 46.68M pieces of content (93.94% of the total), and for Instagram 75.11M (98.44% of the total). These proportions of automatically removed content are consistent with their SoR submissions only if there is no distinction between fully and partially automated decision. LinkedIn’s transparency report covers moderation actions occurred between 25 August and 30 September. It distinguishes between job posts and other content, with the former having a more stringent moderation policy. In its report, LinkedIn claimed to have removed all kinds of violating content detected automatically 99% of the time, whereas for submitted SoRs it was 86% —a relatively small discrepancy.
@@
TikTok, YouTube, Snapchat. TikTok’s transparency report lacks several details and their actions data only covers the month of September 2023, shortcomings that are acknowledged therein. Based on the more detailed actions for removed content, we found a considerable automation discrepancy. The report specifies that 45% of non-ad content was removed automatically, whereas in the DSA-TDB it was 95%, more than double for a partially overlapping period. In the case of YouTube, even if it is the only platform to use all combinations of automated means without one being predominant —in accordance with its transparency report— we found an intriguing anomaly during our exploratory analysis. In the DSA-TDB data we found a remarkable spike in SoRs applied on a single day —5 October— reaching 732.4K. For reference, YouTube’s daily median is 112.2K. Almost all of these are ads that were disabled because deemed outside the scope of the platform, but via non-automated means, which seems implausible given the very large amount of moderation actions undertaken. Unfortunately, with the SoR and transparency report data available we could not delve into this anomaly in more detail. The DSA transparency report by Snapchat is somewhat peculiar, as they already had in place a global transparency report, released twice a year. Hence, they only integrated this report via a document with some high-level information specific to the EU, in order to comply with the first mandatory DSA reporting period. Nevertheless, we noticed an odd inconsistency. DSA transparency reports must include data about Trusted Flaggers: legal entities responsible for notifying platforms of illegal content. As per Article 19(2) of the DSA, Trusted Flaggers are officially appointed by certain EU bodies. However, at the time of writing no Trusted Flagger has yet been designated. For this reason, no platform reported in their SoRs to have received notifications from Trusted Flaggers. With the exception of Snapchat, for which we found a few hundreds SoRs with notification source type set as Trusted Flagger. Upon further inspection we realized that Snapchat has a similar namesake internal program, [^5] https://values.snap.com/safety/safety-center unrelated to the DSA’s. This homonymy probably caused some confusion in Snapchat’s reporting to the DSA-TDB, where they referred to their internal program instead of the intended one.
@@
X. Finally, the transparency report by X includes very detailed data regarding its moderation actions for the period 28 August – 20 October (54 days), with 56K actions being taken on content (36.5% via automated means) and 2M actions on accounts (74.3% via automated means). Yet, all of their SoRs submitted to the DSA-TDB are marked as not using any automated means. In addition, during the 60-day period of our study —which partially overlaps the transparency report’s period— X only submitted 264K SoRs, compared to 2.05M moderation actions declared in the transparency report. These actions covered 14 violation categories, which are similar but not the same as the DSA-TDB categories. For content, violent speech and abuse & harrassment were the most common: 25.5K (47%) and 12.1K (22%), respectively. For accounts, platform manipulation & spam and child sexual exploitation: 1.89M (95%) and 60K (3%) violations, respectively. However, 99.8% of all SoRs submitted by X to the DSA-TDB targeted synthetic media, mostly pornography or sexualized content, as shown in Figures 3 and 1, which appears to be inconsistent with X’s transparency report. Moreover, all of X’s SoRs in the DSA-TDB reported a decision application delay of zero days, without any use of automation, which again seems implausible.
Discussion

### 20 Jan 2024 v2

@@ 4.1. Grounds for Decision
Concerning the decisions grounds, for almost every platform incompatible content was the overwhelming majority (more than 99%), up to all instances in the case of LinkedIn. The exception was X, which indicated illegal content for all of its SoRs. Curiously, only Snapchat indicated an overlap of incompatible and illegal content in 12.2% of their SoRs. None of the other social media platforms used this attribute.
Figure 2. Platform-wise distribution of infringement categories of statement of reasons (SoR), seriated using Manhattan distance across platforms and categories, so that similar platforms are positioned close to one another.
@@
Curiously, category addition was merely used by Pinterest and X, but in less than 0.04% of their SoRs. As for keywords, only half of the social media platforms used them for category specification: Pinterest, Snapchat, X, and YouTube. However, these are not always informative. For instance, circa 83% of YouTube’s SoRs have a keyword —all of them other— which YouTube explained by specifying that “[their] mapping has not yet been completed”. Only 11% of Snapchat’s SoRs have a keyword, 87% of these being sexual abuse material and the rest mainly online bullying intimidation. About 88% of Pinterest’s SoRs have a keyword (with a few having two), predominantly adult sexual material. Interestingly, all of X’s SoRs have a keyword, with a few having two or three. In total X used 17 distinct keywords, more than any other social media. However, in almost half (43%) of its SoRs the keyword is other (detailed as a policy violation), while in 49% it is adult sexual material. In the remaining SoRs the keywords are primarily child sexual abuse material, nudity, and goods/services not permitted.
Table 2. Platform-wise distribution of decision types, ordered alphabetically and by type mean across platforms.
@@
Facebook $42.87\%$ $57.13\%$
Instagram $44.11\%$ $55.89\%$
LinkedIn $99.77\%$ $0.23\%$
Pinterest $99.95\%$ $>0\%$ $0.04\%$
Snapchat $80.15\%$ $19.85\%$
TikTok $98.30\%$ $1.15\%$ $0.55\%$
X $71.44\%$ $6.58\%$ $21.98\%$
YouTube $98.04\%$ $1.96\%$
Type mean $79.33\%$ $17.58\%$ $3.09\%$
Figure 3. Platform-wise distribution for decisions of the visibility type, seriated using Manhattan distance across platforms and sub-types, so that similar platforms are positioned close to one another.
4.2. Types of Restriction Decisions
There are four main types of restriction decisions: account (suspension, termination), monetary (suspension, termination, other), provision (partial/total suspension/termination of the service), and visibility (restriction for the content). Platforms must indicate a single sub-type of one of the four types, but these types are not mutually exclusive. In other words, a SoR can contain multiple types of restrictions, with at least one of them being mandatory. Nonetheless, all of the social media platforms only set a single restriction type. As detailed in Table 2, most were in visibility (79.3% across platforms), followed by account (17.6%) and provision (3.1%). Surprisingly, there was not a single SoR with a monetary restriction decision. For visibility sub-types, shown in Figure 3, we noticed a similar platform grouping to that of infringement categories:
  1. (1)
@@ 4.3. Content Type
The required attribute content type is a list with one or more of the following predefined values: app, audio, image, product, synthetic media (e.g., content generated via AI), text, video, and other. Despite the possibility to indicate multiple values, all eight platforms only specified a single value for all of their SoRs, except YouTube, which however indicated multiple values only in a handful of cases ($<$0.01%). We found that social media platforms greatly differ in how they specify the content type of their SoRs, as shown on the left side of Figure 4. Pinterest and LinkedIn only set content type other, and did not map any of their internal types to one or more of the predefined values in the DSA-TDB. YouTube also had a remarkable share of other content (93.9%), prevalently ads. Facebook and Instagram have similar distributions, with most content being other (more than 55%), all detailed as accounts, which is consistent with their distributions of decision types in Table 2. Incidentally, Facebook was the only social media with content type product (1.8%). More than half of Snapchat’s moderated content was video, with other content (24%) being mainly accounts. Still, 16.6% of other was labelled as multi-media, instead of using multiple predefined values as prescribed by the DSA-TDB. Interestingly, TikTok’s moderated content was mostly text (55.5%) and not video (33.5%), with a small share of other content (1.3%), primarily accounts. Unexpectedly, moderated content by X was overwhelmingly synthetic media (99.8%), with the second being audio (0.01%) —the only social media to use this type— and only a handful of text.
Figure 4. Distribution of moderated content type by platform. For the detailed other content type (right side of the figure), labels were simplified for space reasons.
4.4. Timeliness
@@
Artice 17 of the DSA requires that platforms communicate their SoRs without undue delay to allow close to real-time updates. Pinterest, Snapchat, and TikTok started submitting SoRs on the very day that the DSA-TDB was made available —that is, on 25 September 2023. Facebook, Instagram, YouTube, and LinkedIn the following day. X was last in line and started a week later, on 3 October 2023. As sketched in Figure 5, we then analyzed the delay with which each social media communicated their SoRs, measured as the difference (in days) between the date of application of the moderation intervention and the date of creation of the corresponding SoR. In addition, a prompt response by platforms in moderating an infringing content is arguably desirable. Hence, we also measured the application delay (in days) of the moderation decision with respect to the creation of the infringing content.
Figure 5. We analyzed the timeliness of submitted moderation actions based on: lag between the creation of the infringing content and the application of the moderation decision (application delay); and lag between this application and the communication to the DSA-TDB, which implies the creation of a statement of reasons (communication delay). Labels in monospaced font refer to the database fields used for computing the delays.
@@
Regarding communication delay, LinkedIn, Pinterest, and Snapchat communicated their SoRs on the same day of application with little to no fluctuations, as shown in Figure 6. Despite the late start, X caught up on the first week of submissions by communicating content moderated since 25 September. Afterwards, it began to communicate SoRs the same day of their application. On the contrary, TikTok started by stably submitting on the same day of application, but in the second half of the 100-day period it communicated SoRs with an uprising mean delay of up to eight days. Facebook and Instagram communicated their SoRs with a mean delay of one day, with a few delay surges common to both Meta platforms. On the other hand, YouTube had a minimum delay of one day, but with marked fluctuations throughout the period of analysis.
Figure 6. Platform-wise daily mean of the delay (measured in days) in the communication of statements of reasons to the DSA-DTB.
@@
Concerning application delay, we observed great heterogeneity in its distribution among platforms. For this reason, in their analysis and visualization we opted for using few categorical cuts, as shown in Figure 7. In addition to visualizing the distributions of applications delay, we also report in the figure the 80% winsorized mean (W) as a robust measure of centrality that lessens the influence of outliers. Oddly, X was the only platform that always applied a decision on the same day the infringing content was created (W$=$0). TikTok had the second lowest delay, with 89% of decisions being applied the day of content creation. Next, LinkedIn applied its decision within 7 days in 89.8% of their SoRs, while Instagram in 81.7%. Snapchat had a similar value for the same timeframe (81.4%), but in 14.1% of SoRs it took more than 30 days to apply its decision. Facebook presented a similar but more marked situation, with 72% and 21.5%, respectively. YouTube, on the other hand, had a substantial amount of decisions that took more than 30 days (70.7%), while also having many within 15–30 days (17.5%). However, the distribution with the highest mean and dispersion was that of Pinterest, as 17.5% of decisions were applied the same day of content creation while 70.8% after 30 days, with many even taking hundreds of days.
Figure 7. Distribution of decision application delays (with respect to the date of content creation) by social media platform, ordered by their 80% winsorized means (W).
4.5. Automated Means
@@
Automation can be applied at various steps in the content moderation process. It is often used to scale up the detection of offending content and to speed up the decision on how to moderate it. Consequently, the DSA-TDB includes two attributes to describe the possible use of automation for moderation: automated detection (yes, no) and automated decision (not, partially, fully). Figure 8 provides a combined view of the use of automation for detection and decision by social media platforms. Again, we see different approaches being adopted. Surprisingly, none of the SoRs submitted by X specified the use of automated means whatsoever. Snapchat is second in terms of non-automated means in their SoRs (74%). Pinterest mainly uses non-automated detection with partially automated decision (98%), whereas LinkedIn mainly uses automated detection together with non-automated decision (64%). YouTube is the only platform that uses all combinations of automated means, with automated detection and non-automated decision being the most common (51%), followed by fully automated means (28%), and non-automated means (16%). Once more, Instagram and Facebook have similar distributions, with automated detection and partially automated decision being the most common for both, with 93% and 99%, respectively. Almost opposite of X, TikTok predominantly indicates fully automated means for moderation (92%).
Figure 8. Platform-wise distribution of automated infringement detection and decision, ordered by an ad hoc automation index.
@@
To provide a holistic view of social media behavior with respect to automated moderation, we converted the boolean and ordinal attributes of the DSA-TDB into two numeric indices of automated detection and automated decision. Both indices are defined in the [0, 1] range and are valued, for each platform, depending on the frequency with which automated detection and decision occurred in the platform’s SoRs. We were then able to position each platform in the bi-dimensional space depicted in Figure 9, where dot size is proportional to the number of SoRs per MAAR, and dot color represents the mean of the automated detection and automated decision indices. The platform positioning within the newly defined automation space reveals four groups of platforms with overall similar degrees of automation: low (X and Snapchat), modest (Pinterest), substantial (LinkedIn, YouTube, Instagram, and Facebook), and high (TikTok). At the platform level, the ratio in SoRs per MAAR is moderately correlated with an automated index ($r=0.6)$. However, the correlation between number of SoRs per MAAR and automated detection is much lower ($r=0.3$) compared to automated decision ($r=0.9$). In particular, Pinterest stands apart from others platforms due to its very low level of automated detection but very high level of partially automated decision.
Figure 9. Use of automated means by social media platforms for the detection of infringements and decision on these. Size is proportional to the number of statements of reasons (SoRs) per monthly authenticated active recipients (MAAR). In general, the bigger this ratio is, the higher the automated index ($r=0.6)$, with Pinterest being somewhat of an outlier due to its marginal automated detection.
4.6. Comparison with Transparency Reports
@@
Within the DSA framework there are multiple mechanisms to ensure transparency in content moderation, including the DSA-TDB that we analyzed so far, and the DSA Transparency Reports. Both resources cover similar aspects of the content moderation process, albeit at different granularities. As such, there is overlapping between the information contained in the two resources, which opens up the possibility to carry out thorough platform-wise comparisons. In this section we leverage this opportunity to carry out the first systematic comparison between the data submitted by each social media to the DSA-TDB and that declared in the corresponding transparency report. In detail, for each platform and for each dimension of moderation (e.g., timeliness, automation, etc.), we look for the corresponding information in that platform’s most recent DSA Transparency Report. We then compare the two figures to corroborate or challenge the self-declared information provided by each platform. This comparison serves as a validation step of the DSA-TDB data, and allows for spotting inconsistencies and discrepancies. However, in general, we were not able to compare each dimension of moderation for each of the eight platforms, due to differences in the data they provided in their transparency reports. Still, all platforms —with the exception of Snapchat— provided detailed information about the use of automated means for moderation, for which we found many inconsistencies. Further, in the case of X, we found inconsistencies across multiple aspects. In the following, we present the results of our comparisons, starting from the platforms for which we found less issues.
@@
Pinterest, Facebook, Instagram, LinkedIn. Pinterest’s self-reporting was overall congruous and detailed. Interestingly, it was the only social media to use three automation levels in their transparency report —manual, hybrid, and automated— similarly to the organization of the DSA-TDB. Unsurprisingly, Facebook and Instagram share similar transparency reports, based on Meta guidelines. In their reports, both state the heavy use of automated means, including full automation for both detection and decision. However, none of their SoRs specified fully automated decisions, let alone full automation overall, as shown in Figure 8. For the reporting period 25 April – 30 September 2023 (159 days), Meta platforms automatically removed the following organic content (i.e., non-paid): for Facebook, 46.68M pieces of content (93.94% of the total), and for Instagram 75.11M (98.44% of the total). These proportions of automatically removed content are consistent with their SoR submissions only if there is no distinction between fully and partially automated decision. LinkedIn’s transparency report covers moderation actions occurred between 25 August and 30 September. It distinguishes between job posts and other content, with the former having a more stringent moderation policy. In its report, LinkedIn claimed to have removed all kinds of violating content detected automatically 99% of the time, whereas for submitted SoRs it was 85% —a relatively small discrepancy.
@@
TikTok, YouTube, Snapchat. TikTok’s transparency report lacks several details and their data only covers the month of September 2023, shortcomings that are acknowledged therein. Based on the more detailed actions for removed content, we found a considerable automation discrepancy. The report specifies that 45% of non-ad content was removed automatically, whereas in the DSA-TDB it was 95%, more than double for a partially overlapping period. In the case of YouTube, even if it is the only platform to use all combinations of automated means without one being predominant —in accordance with its transparency report— we found intriguing anomalies. In the DSA-TDB we saw a remarkable spike in SoRs applied on 5 October 2023, reaching 732.4K. For reference, YouTube’s daily median is 112.8K. Almost all of these are ads disabled because deemed outside the scope of the platform, but via non-automated means, which seems implausible. Furthermore, there were three similar spikes in mid December with even higher single-day volumes (¿1.25M each) corresponding to a majority of SoRs having automated detection but non-automated decision. Unfortunately, with the data available we could not delve deeper into these anomalies. Snapchat’s transparency report is somewhat peculiar, as they already had in place a namesake global transparency report, released twice a year. Hence, they only integrated this report via a document with some high-level information specific to the EU, in order to comply with the first mandatory DSA reporting period. Nevertheless, we noticed an odd inconsistency. DSA transparency reports must include data about Trusted Flaggers: legal entities responsible for notifying platforms of illegal content (European Parliament and Council, 2022). As per Article 19(2) of the DSA, Trusted Flaggers are officially appointed by certain EU bodies. However, at the time of writing no Trusted Flagger has yet been designated. For this reason, no platform reported in their SoRs to have received notifications from Trusted Flaggers, with the exception of Snapchat, for which we found a few hundreds SoRs with notification source type set as Trusted Flagger. Upon further inspection we realized that Snapchat has a similar namesake internal program, [^8] https://values.snap.com/safety/safety-center unrelated to the DSA’s one. This homonymy probably caused some confusion in Snapchat’s reporting to the DSA-TDB, where they referred to their internal program instead of the intended one.
@@
X. Finally, the transparency report by X includes very detailed data regarding its moderation actions for the period 28 August – 20 October 2023 (54 days), with 56K actions being taken on content (36.5% via automated means) and 2M actions on accounts (74.3% via automated means). Yet, all of their SoRs submitted to the DSA-TDB are marked as not using any automated means. In addition, during the 100-day period of our study —which partially overlaps with the transparency report’s period— X only submitted 466.4K SoRs, compared to 2.05M moderation actions declared in the transparency report. Those actions covered 14 violation categories, which are similar but not the same as the DSA-TDB categories. For content, violent speech and abuse & harrassment were the most common: 25.5K (47%) and 12.1K (22%), respectively. For accounts, platform manipulation & spam and child sexual exploitation: 1.89M (95%) and 60K (3%) violations, respectively. However, 99.8% of all SoRs submitted by X to the DSA-TDB targeted synthetic media, mostly pornography or sexualized content, as shown in Figures 4 and 2, which appears to be inconsistent with X’s transparency report. Moreover, all of X’s SoRs in the DSA-TDB reported a decision application delay of zero days, without any use of automation, which again, seems very implausible.
5. Discussion